OpenStudy (anonymous):

How many grams of solute are present in 11.0 ml of .786 M K2Cr2O7

OpenStudy (abb0t):

Well, first of all you have 11.0 mL, which you need to convert to Liters. Then, you mutliply it by 0.786 \(\sf \frac{mol}{L}\) of potassium chromate. and finally multiply by the molar mass to get grams.

OpenStudy (anonymous):

so whats the answer?

OpenStudy (abb0t):

You tell me.

OpenStudy (abb0t):

I told you what to do already. Molar mass of potassium dichromate is: 294.185 \(\sf \frac{g}{mol}\)

OpenStudy (anonymous):

2.54352 is this right?

OpenStudy (ipwnbunnies):

dichromate* :3

OpenStudy (abb0t):

Yes.

OpenStudy (anonymous):

If 13.7g of (NH4)2SO4 is dissolved in enough water to form 230mL of solution, what is the molarity of the solution?

OpenStudy (abb0t):

No.

OpenStudy (anonymous):

no, what steps do i take

OpenStudy (abb0t):

You find the number of moles, then divide by 230 mL \(\times\) 10\(^3\) to get your molarity

OpenStudy (abb0t):

You should have units: \(\sf \color{red}{\frac{mol}{L}}\)

OpenStudy (anonymous):

how many moles?

OpenStudy (anonymous):

Hey abb0t could you please help me undertsand my question ? V 1.1 x 103 mm Hg 1.4 x 103 mm Hg 1.5 x 102 mm Hg 3.9 x 102 mm Hg

OpenStudy (anonymous):

I don't know what steps to take

OpenStudy (abb0t):

You need to find the number of moles by multiplying by the molar mass. You can use the periodic table to find that. If you do not have one on hand, you can go here: http://ptable.com/

OpenStudy (anonymous):

this was the question you were helping me with Initially, a 1.50-liter compressible container of 0.35 moles of a gas exerts a pressure of 740 millimeters of mercury at a temperature of 275 Kelvin. What is the pressure when the container is compressed to 0.50 liters, the moles of gas reduces to 0.20 moles, and the temperature changes to 32 degrees Celsius? Answer 1.1 x 103 mm Hg 1.4 x 103 mm Hg 1.5 x 102 mm Hg 3.9 x 102 mm Hg

OpenStudy (anonymous):

is this right .0024978?

OpenStudy (anonymous):

YOURE WRONG

OpenStudy (abb0t):

divide by the volume!!!! Molarity is moles per liter!!!